OpenStudy (sette):

The length of a rectangle is 3 more than twice its width. The perimeter is 48 feet. Find the width. 16 7 8 21 24 2L+2W=48 L=2W+3

OpenStudy (anonymous):

L = 3 + 2W So, Perimeter = 2L + 2W 48 = 2(3 + 2W) + 2W 48 = 6 + 4W + 2W 6W = 42 Divide by 6 and find W from here..

OpenStudy (sette):

8

OpenStudy (anonymous):

42 divided by 6 gives you 8???

OpenStudy (sette):

by bad , 7 .

OpenStudy (anonymous):

\(6 \times 7 = 42\)

OpenStudy (anonymous):

Yes the Width is 7 feet..

OpenStudy (sette):

Thank You .